# Framing global structural identifiability in terms of parameter symmetries

### v1

@@ 2.1 Global structural identifiability and the standard differential algebra approach
$\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\boldsymbol{\theta}^{\star})\quad\forall{t}\in\mathbb{R}\implies{\theta}_{j}=\theta_{j}^{\star}.$ (3)
The idea is essentially that a parameter is structurally identifiable if a change in the parameter results in a change in the output. In the case when the non-linear functions $\mathbf{f}$ and $\mathbf{h}$ defining the mechanistic model in Eq. (1) are rational functions of the independent and dependent variables, the standard differential algebra approach for conducting a structural identifiability analysis [16] entails conducting four steps (Algorithm 1). The intuition behind the standard differential algebra approach, which re-writes the original first-order system of ODEs as an input-output system as in Eq. (2), is that these two systems are equivalent with respect to outputs as the latter system constitutes an exhaustive summary of the model of interest [17]. Technically, this implies that by generating outputs $\mathbf{y}$ from a particular solution of the original first-order system using specific parameters $\boldsymbol{\theta}$, the same outputs $\mathbf{y}$ also solve the input-output system [17]. Furthermore, the parameter combinations realised in the input-output system define the observed outputs $\mathbf{y}$ owing to the uniqueness of solutions of systems of ODEs. Next, we consider an alternative approach for analysing structural identifiability based on so-called Lie symmetries.
Input: A system of first-order ODEs with associated observed outputs as in Eq. (1) where the functions $\mathbf{f}$ and $\mathbf{h}$ are rational functions of the states and outputs.
Output: The globally structurally identifiable parameter quantities.
Step 1: Output equations. Rewrite the model equations in Eq. (1) as input-output equations as in Eq. (2);
Step 2: Monic polynomial equations. Write input-output equations as monic polynomial equations;
Step 3: Coefficients. Extract polynomial coefficients;
Step 4: Create injective parameter map. Consider the map from parameter space to the polynomial coefficients. Find “identifiable parameter combinations” for which the map is one-to-one and return these.
Algorithm 1 The standard differential algebra approach for conducting a structural identifiability analysis.
2.2 Extending classical Lie symmetries to full symmetries acting on parameters
We consider solution curves of the first-order ODE system in Eq. (1) describing their dependent variables as functions of the independent variable and the rate parameters. Furthermore, we introduce the parameter $\varepsilon$ which parameterises such solution curves. Then, we are interested in a (one-parameter) Lie transformation $\Gamma_{\varepsilon}:\mathbf{R}\times\mathbf{R}^{p}\times\mathbf{R}^{n}\mapsto\mathbf{R}\times\mathbf{R}^{p}\times\mathbf{R}^{n}$ which maps a solution curve to another solution curve according to
$\Gamma_{\varepsilon}:\left(t,\boldsymbol{\theta},\mathbf{x}\right)\mapsto\left(\hat{t}(t,\boldsymbol{\theta},\mathbf{x},\varepsilon),\hat{\boldsymbol{\theta}}(t,\boldsymbol{\theta},\mathbf{x},\varepsilon),\hat{\mathbf{x}}(t,\boldsymbol{\theta},\mathbf{x},\varepsilon)\right),$ (4)
where we have used the hat notation to denote the transformed coordinates. In particular, the trivial transformation is defined by setting $\varepsilon=0$, i.e. $\hat{t}(\varepsilon=0)=t$, $\hat{\mathbf{x}}(\varepsilon=0)=\mathbf{x}$ and $\hat{\boldsymbol{\theta}}(\varepsilon=0)=\boldsymbol{\theta}$. Importantly, the transformed coordinates are continuous functions of $\left(t,\boldsymbol{\theta},\mathbf{x}\right)$ corresponding to the independent and dependent variables as well as the parameters, and they are parameterised by the transformation parameter $\varepsilon$. More precisely, the Lie transformation $\Gamma_{\varepsilon}$ is a $\mathcal{C}^{\infty}(\mathbf{R}\times\mathbf{R}^{p}\times\mathbf{R}^{n})$ diffeomorphism [12, 13, 14, 15] which implies that we can Taylor expand each of the transformed coordinates around $\varepsilon\approx 0$ as follows
$\begin{split}\hat{t}(t,\boldsymbol{\theta},\mathbf{x},\epsilon)&=t+\left.\frac{\mathrm{d}\hat{t}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\varepsilon+\mathcal{O}\left(\varepsilon^{2}\right)=t+\xi(t,\mathbf{\theta},\mathbf{x})\varepsilon+\mathcal{O}\left(\varepsilon^{2}\right),\\ \hat{\theta}_{\ell}(t,\boldsymbol{\theta},\mathbf{x},\epsilon)&=\theta_{\ell}+\left.\frac{\mathrm{d}\hat{\theta_{\ell}}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\varepsilon+\mathcal{O}\left(\varepsilon^{2}\right)=\theta_{\ell}+\chi_{\ell}(t,\mathbf{\theta},\mathbf{x})\varepsilon+\mathcal{O}\left(\varepsilon^{2}\right),\quad{\ell\in\left\{1,\ldots,p\right\}},\\ \hat{x}_{i}(t,\boldsymbol{\theta},\mathbf{x},\epsilon)&=x_{i}+\left.\frac{\mathrm{d}\hat{x_{i}}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\varepsilon+\mathcal{O}\left(\varepsilon^{2}\right)=x_{i}+\eta_{i}(t,\mathbf{\theta},\mathbf{x})\varepsilon+\mathcal{O}\left(\varepsilon^{2}\right),\quad{i\in\left\{1,\ldots,n\right\}}.\end{split}$ (5)
Here, we refer to the unknown functions denoted by $\xi$, $\eta_{i}$ and $\chi_{j}$ as the infinitesimals [12]. A convenient piece of notation is to introduce the vector field known as the infinitesimal generator of the Lie group $X$ given by
$X=\xi(t,\boldsymbol{\theta},\mathbf{x})\partial_{t}+\sum_{i=1}^{n}\eta_{i}(t,\boldsymbol{\theta},\mathbf{x})\partial_{x_{i}}+\sum_{\ell=1}^{p}\chi_{\ell}(t,\boldsymbol{\theta},\mathbf{x})\partial_{\theta_{j}},$ (6)
which corresponds to the infinitesimal description of the Lie transformation $\Gamma_{\varepsilon}$ in Eq. (4). Importantly, classical symmetries acting on dependent and independent variables correspond to full symmetries for which the parameter infinitesimals are all zero, i.e. $\chi_{1}=\ldots=\chi_{p}=0$. Using the infinitesimals, we generate the symmetry itself by solving the following ODE system [12]:
$\begin{split}\dfrac{\mathrm{d}\hat{t}}{\mathrm{d}\varepsilon}&=\xi(\hat{t},\hat{\boldsymbol{\theta}},\hat{\mathbf{x}}),\quad\hat{t}(\varepsilon=0)=t,\\ \dfrac{\mathrm{d}\hat{x_{i}}}{\mathrm{d}\varepsilon}&=\eta_{i}(\hat{t},\hat{\boldsymbol{\theta}},\hat{\mathbf{x}}),\quad\hat{x_{i}}(\varepsilon=0)=x_{i},\quad{i\in\left\{1,\ldots,n\right\}},\\ \dfrac{\mathrm{d}\hat{\theta_{\ell}}}{\mathrm{d}\varepsilon}&=\chi_{\ell}(\hat{t},\hat{\boldsymbol{\theta}},\hat{\mathbf{x}}),\quad\hat{\theta_{\ell}}(\varepsilon=0)=\theta_{\ell},\quad{\ell\in\left\{1,\ldots,p\right\}},\\ \end{split}$ (7)
and hence the symmetry $\Gamma_{\varepsilon}$ is completely characterised by its infinitesimals and its generating vector field $X$. To find these infinitesimals, we need to extend these Lie transformations slightly.
2.2.1 The symmetry conditions
To characterise a Lie transformation $\Gamma_{\varepsilon}$ as a symmetry of a system of differential equations, we must represent the system of ODEs in Eq. (1) as a geometrical object on which certain functions vanish. Then Lie transformations maps solutions to other solutions in such a way that this geometrical object is invariant under transformations [14]. Given $n$ states $\mathbf{x}\in\mathbb{R}^{n}$ depending on $t\in\mathbb{R}$ and $\boldsymbol{\theta}\in\mathbb{R}^{p}$, we have $n$ derivatives $\dot{\mathbf{x}}\in\mathbb{R}^{n}$ and thus $(t,\boldsymbol{\theta},\mathbf{x},\dot{\mathbf{x}})$ constitutes a point in the so-called first Jet space [14] $\mathcal{J}^{(1)}=\mathbb{R}\times\mathbb{R}^{p}\times\mathbb{R}^{n}\times\mathbb{R}^{n}$. Moreover, the solution manifold $\mathcal{S}$ given by
$\mathcal{S}\coloneqq\left\{(t,\boldsymbol{\theta},\mathbf{x},\dot{\mathbf{x}})\in\mathcal{J}^{(1)}:\Delta_{1}(t,\boldsymbol{\theta},\mathbf{x},\dot{\mathbf{x}})=\dot{\mathbf{x}}-\mathbf{f}\left(t,\mathbf{x},\boldsymbol{\theta}\right)=\mathbf{0}\right\},$ (8)
is a subvariety of the Jet space $\mathcal{J}^{(1)}$ characterised by the vanishing of the function $\Delta_{1}$ [14] and symmetries preserve this subvariety. Also, given an index $j\in\{1,\ldots,n\}$, each derivative $\dot{x}_{j}(t,\boldsymbol{\theta})$ is uniquely defined by the original state $x_{j}(t,\boldsymbol{\theta})$. Thus, any solution curve $\mathbf{x}(t,\boldsymbol{\theta})=\left(x_{1}(t,\boldsymbol{\theta}),\dots,x_{n}(t,\boldsymbol{\theta})\right)$ has a uniquely defined extension in terms of an induced function $\mathrm{pr}^{(1)}\mathbf{x}(t):\mathbb{R}^{n}\mapsto\mathbb{R}^{n}\times\mathbb{R}^{n}$ referred to as the first prolongation [14] which is defined by
$\mathrm{pr}^{(1)}\mathbf{x}(t,\boldsymbol{\theta})=(\mathbf{x}(t,\boldsymbol{\theta}),\dot{\mathbf{x}}(t,\boldsymbol{\theta}))=\left(x_{1}(t,\boldsymbol{\theta}),\dots,x_{n}(t,\boldsymbol{\theta}),\dot{x}_{1}(t,\boldsymbol{\theta}),\ldots,\dot{x}_{n}(t,\boldsymbol{\theta})\right).$ (9)
Here, we have included first derivatives of the states in the original solution curve. Moreover, by introducing the function referred to as the total derivative $D_{t}$ [13] defined by
$D_{t}=\partial_{t}+\sum_{i=1}^{n}\dot{x_{i}}\partial_{x_{i}},$ (10)
we can define the notion of the first prolongation of a transformed solution curve
$\begin{split}\mathrm{pr}^{(1)}\hat{\mathbf{x}}(t,\boldsymbol{\theta},\mathbf{x},\varepsilon)&=\left(\hat{\mathbf{x}}(t,\boldsymbol{\theta},\mathbf{x},\varepsilon),\hat{\dot{\mathbf{x}}}(t,\boldsymbol{\theta},\mathbf{x},\dot{\mathbf{x}},\varepsilon)\right)=\left(\hat{x}_{1}(\varepsilon),\dots,\hat{x}_{n}(\varepsilon),\frac{D_{t}\hat{x}_{1}(\varepsilon)}{D_{t}\hat{t}(\varepsilon)},\ldots,\frac{D_{t}\hat{x}_{n}(\varepsilon)}{D_{t}\hat{t}(\varepsilon)}\right),\\ \end{split}$ (11)
where transformed derivatives $\hat{\dot{\mathbf{x}}}(\varepsilon)$ are defined in terms of the total derivative $D_{t}$. This, in turn, enables us to introduce the first prolongation of a Lie transformation, denoted by $\Gamma_{\varepsilon}^{(1)}:\mathcal{J}^{(1)}\mapsto\mathcal{J}^{(1)}$, which is given by
$\Gamma_{\varepsilon}^{(1)}:\left(t,\boldsymbol{\theta},\mathrm{pr}^{(1)}\mathbf{x}(t,\boldsymbol{\theta})\right)\mapsto\left(\hat{t}(t,\boldsymbol{\theta},\mathbf{x},\varepsilon),\hat{\boldsymbol{\theta}}(t,\boldsymbol{\theta},\mathbf{x},\varepsilon),\mathrm{pr}^{(1)}\hat{\mathbf{x}}(t,\boldsymbol{\theta},\mathbf{x},\varepsilon)\right).$ (12)
Importantly, for any Lie transformation $\Gamma_{\varepsilon}$ its first prolongation $\Gamma_{\varepsilon}^{(1)}$ is uniquely defined. Provided first prolongations of Lie transformations, it is straightforward to characterise a Lie transformation $\Gamma_{\varepsilon}$ as a symmetry of the ODEs in Eq. (1) using its first prolongation $\Gamma_{\varepsilon}^{(1)}$. Since we want $\Gamma_{\varepsilon}$ to map a solution curve to another solution curve, $\Gamma_{\varepsilon}$ is a symmetry if and only if its first prolongation $\Gamma_{\varepsilon}^{(1)}$ preserves the solution manifold according to $\Gamma_{\varepsilon}^{(1)}:\mathcal{S}\mapsto\mathcal{S}$ implying that it satisfies the following symmetry conditions:
$\begin{split}\hat{\dot{\mathbf{x}}}(\varepsilon)&=\mathbf{f}\left(\hat{t}(\varepsilon),\hat{\boldsymbol{\theta}}(\varepsilon),\hat{\mathbf{x}}(\varepsilon)\right)\\ \hat{\mathbf{y}}(\varepsilon)&=\mathbf{h}\left(t,\boldsymbol{\theta},\mathbf{x}\right)\end{split}\quad\textrm{whenever}\quad\begin{split}\dot{\mathbf{x}}&=\mathbf{f}\left(t,\boldsymbol{\theta},\mathbf{x}\right)\\ \mathbf{y}&=\mathbf{h}\left(t,\boldsymbol{\theta},\mathbf{x}\right)\end{split}.$ (13)
Essentially these conditions state that if we start with a prolonged solution curve in Eq. (9), the prolonged transformed curve in Eq. (11) should also be solution of the ODE system in Eq. (1). In other words, a symmetry $\Gamma_{\varepsilon}$ of the ODE system Eq. (1) maps solution curves to solution curves. Furthermore, in light of the definition of global structural identifiability of parameters (Defn. 3), transformations by full symmetries leave the outputs $\mathbf{y}$ invariant [6, 7, 8, 9, 10] implying that $\hat{\mathbf{y}}(\varepsilon)=\mathbf{y}~{}\forall\varepsilon\in\mathbb{R}$.
2.2.2 The linearised symmetry conditions
Typically, we do not work with the symmetry conditions themselves but instead we use the equivalent infinitesimal descriptions. Just like the infinitesimal description of $\Gamma_{\varepsilon}$ is given by $X$, the infinitesimal description of $\Gamma_{\varepsilon}^{(1)}$ is given by the vector field known as the first prolongation of the infinitesimal generator of the Lie group $X^{(1)}$. This vector field is defined by
$X^{(1)}=X+\sum_{i=1}^{n}\eta_{i}^{(1)}\partial_{\dot{x_{i}}},\quad{i\in\left\{1,\ldots,n\right\}},$ (14)
where the first prolongations of the infinitesimals are calculated by using the prolongation formula [13]
@@ 2.2.2 The linearised symmetry conditions
$\eta^{(1)}_{i}\left(t,\boldsymbol{\theta},\mathbf{x},\dot{x}_{i}\right)=D_{t}\eta_{i}\left(t,\boldsymbol{\theta},\mathbf{x}\right)-\dot{x_{i}}\xi\left(t,\boldsymbol{\theta},\mathbf{x}\right),\quad{i\in\left\{1,\ldots,n\right\}}.$ (15)
The prolongation formula can be used recursively in order to allow us to account for higher order derivatives, as well implying that Lie symmetries can be used to analyse systems of higher order ODEs in addition to first-order systems. Given the first prolongation $X^{(1)}$, the infinitesimal descriptions of the symmetry conditions are known as the linearised symmetry conditions which are given by
$\begin{split}X^{(1)}\left(\dot{\mathbf{x}}-\mathbf{f}\left(t,\mathbf{x},\boldsymbol{\theta}\right)\right)&=\mathbf{0}\\ X\left(\mathbf{y}\right)&=\mathbf{0}\end{split}\quad\textrm{whenever}\quad\begin{split}\dot{\mathbf{x}}&=\mathbf{f}\left(t,\mathbf{x},\boldsymbol{\theta}\right)\\ \mathbf{y}&=\mathbf{h}\left(t,\mathbf{x},\boldsymbol{\theta}\right)\end{split}.$ (16)
Essentially, the linearised symmetry conditions in Eq. (16) correspond to the $\mathcal{O}(\varepsilon)$ terms in the Taylor expansions of the symmetry conditions in Eq. (13). Importantly, these linearised symmetry conditions state that the outputs $\mathbf{y}=\mathbf{h}\left(t,\boldsymbol{\theta},\mathbf{x}\right)$ are so-called zeroth order differential invariants of $X$, and these are referred to as canonical coordinates [12, 13] in the literature on classical Lie symmetries.
3 Results
@@
3.1 Parameter symmetries are Lie transformations acting as re-parameterisations that preserve observed outputs
We first define a special type of full symmetry referred to as a parameter symmetry.
Defn. 2 (Parameter symmetries).
@@
Let $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$ be a full symmetry that is restricted to the parameters of the system of output ODEs in Eq. (2) defined by
@@
where the target functions $\hat{\boldsymbol{\theta}}$ depend solely on the parameters $\boldsymbol{\theta}$ in addition to the parameter $\varepsilon$. In other words, the independent variable $t$ and the dependent variables $\mathbf{y}$ are invariant under the action of $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$, implying that for any solution curves the following conservation property holds
@@
$X_{\boldsymbol{\theta}}=\sum_{\ell=1}^{p}\chi_{\ell}(\boldsymbol{\theta})\partial_{\theta_{\ell}}.$ (19)
Then $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$ in Eq. (17) is a parameter symmetry of the system of output ODEs in Eq. (2) if its infinitesimal generator $X_{\boldsymbol{\theta}}$ solves the linearised symmetry conditions given by
@@
In terms of infinitesimals, parameter symmetries are full symmetries characterised by two properties. First, the infinitesimals corresponding to the independent and dependent variables $t$ and $\mathbf{x}$, respectively, are zero, i.e. $\xi=\eta_{1}=\ldots=\eta_{n}=0$. Second, as stated in Defn. 20, the infinitesimals corresponding to the parameters given by $\chi_{1},\ldots,\chi_{p}$ depend solely on the parameters $\boldsymbol{\theta}$ and they do not depend on the independent and dependent variables.
Essentially, a parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ is a re-parametrisation of the model in Eq. (2) which preserves the observed outputs $\mathbf{y}$. Since these parameter symmetries are restricted to the parameters of the model, we will use the notation $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\hat{\boldsymbol{\theta}}(\varepsilon)$ to describe them henceforth. Next, we define the notion of differential invariants of parameter symmetries.
Defn. 3 (Differential invariants of parameter symmetries).
Consider a parameter symmetry $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$ and its corresponding infinitesimal generator $X_{\boldsymbol{\theta}}$. A non-constant function $I=I\left(t,\frac{\mathrm{d}^{N}\mathbf{y}}{\mathrm{d}t^{N}},\frac{\mathrm{d}^{N-1}\mathbf{y}}{\mathrm{d}t^{N-1}},\ldots,\frac{\mathrm{d}\mathbf{y}}{\mathrm{d}t},\mathbf{y},\boldsymbol{\theta}\right)$ is called a differential invariant of $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$ if it satisfies
@@
From this definition, we immediately see that the independent time variable $t$, all outputs $\mathbf{y}$, and their respective derivatives are themselves differential invariants. In addition to the independent and dependent variables, there are other invariants solely depending on the rate parameters, and to distinguish between these two types of invariants we introduce the notion of parameter invariants (Defn. 22).
Defn. 4 (Parameter invariants of parameter symmetries).
@@
Moreover, for a general output system of ODEs there are potentially many possible re-parametrisations, or, differently put, many possible parameter symmetries $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ (Defn. 20). For instance, consider the trivial parameter symmetry $\Gamma_{0}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\boldsymbol{\theta}$ which is common to all possible output systems. For this parameter symmetry, every single parameter is a parameter invariant, and thus most of the parameter invariants of the trivial parameter symmetry are likely not shared with other parameter symmetries of the same model. In general, parameter symmetries have some parameter invariants that are unique to them, and other parameter invariants that are shared with all other parameter symmetries. We refer to this latter type of parameter invariant as universal parameter invariants.
Defn. 5 (Universal parameter invariants of a model).
@@
Importantly, the independent variable $t$, the outputs $\mathbf{y}$, and all derivatives of the outputs are also universal invariants of all parameter symmetries.
Given the system of output ODEs in Eq. (2), we know that the number of parameter invariants for any parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ is an integer in the set $\{0,\ldots,p\}$ where $p$ is the number of parameters. When we have zero parameter invariants the output ODEs completely lack parameters, and in the case of $p$ parameter invariants then each parameter is itself an invariant which is only true for the trivial parameter symmetry $\Gamma_{0}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\boldsymbol{\theta}$. From now on, we restrict ourselves to output ODEs in the form of Eq. (2) that contain parameters, i.e. we exclude the extreme case where we have $p=0$ parameters. In particular, if the parameter $\theta_{\ell}$ for some $\ell\in\{1,\ldots,p\}$ is a parameter invariant, i.e. $I_{\boldsymbol{\theta}}(\boldsymbol{\theta})=\theta_{\ell}$, then Eq. (22) gives
@@
Consequently, a parameter $\theta_{\ell}$ is invariant if its infinitesimal is zero, i.e. $\chi_{\ell}=0$. Moreover, by generating the corresponding transformation $\hat{\theta}_{\ell}(\varepsilon)$ using the infinitesimal $\chi_{\ell}$ which entails solving the following ODE,

### v3

@@ 2.1 Global and local structural identifiability
$\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\boldsymbol{\theta}^{\star})\quad\forall{t}\in\mathbb{R}\implies{\theta}_{j}=\theta_{j}^{\star}\quad\quad\forall\,\boldsymbol{\theta},\boldsymbol{\theta}^{\star}\in\mathbb{R}^{p}.$ (5)
The idea is essentially that a parameter is globally structurally identifiable if any change in the parameter results in a change in the output. A closely related concept is that of local structural identifiability which means that a local change in parameters results in a change in the output. Here, the effect of changing parameters on observed outputs is analysed in a ball of radius $\rho>0$ centred around a given parameter vector $\boldsymbol{\theta}_{C}\in\mathbb{R}^{p}$, and we denote this ball by $\mathcal{B}_{\rho}\left(\boldsymbol{\theta}_{C}\right)\coloneqq\left\{\boldsymbol{\theta}\in\mathbb{R}^{p}:|\boldsymbol{\theta}-\boldsymbol{\theta}_{C}|<\rho\right\}$.
Defn. 2 (Local structural identifiability of parameters).
An individual parameter $\theta_{j}\in\boldsymbol{\theta}$, ${j}\in{1,\ldots,p}$ is locally structurally identifiable if for almost every value $\boldsymbol{\theta}^{\star}$ and almost all initial conditions the following holds:
@@
$\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\boldsymbol{\theta}^{\star})\quad\forall{t}\in\mathbb{R}\implies{\theta}_{j}=\theta_{j}^{\star}\quad\quad\forall\,\boldsymbol{\theta},\boldsymbol{\theta}^{\star}\in\mathcal{B}_{\rho}\left(\boldsymbol{\theta}_{C}\right),$ (6)
for some radius $\rho>0$ and some centre $\boldsymbol{\theta}_{C}\in\mathbb{R}^{p}$.
A globally structurally identifiable parameter is always locally structurally identifiable, but the converse statement is not always true. Similarly, we can also define globally identifiable parameter combinations [17, 18], such as the sum of two parameters, in terms of functions $f:\mathbb{R}^{p}\mapsto{\mathbb{R}}$ of the parameters (Defn. 7)
Defn. 3 (Global structural identifiability of parameter combinations).
The function $f:\mathbb{R}^{p}\mapsto{\mathbb{R}}$, where $f(\boldsymbol{\theta})$ defines a parameter combination, is globally structurally identifiable if for almost every value $\boldsymbol{\theta}^{\star}$ and almost all initial conditions the following holds:
$\mathbf{y}(t,\boldsymbol{\theta})=\mathbf{y}(t,\boldsymbol{\theta}^{\star})\quad\forall{t}\in\mathbb{R}\implies{f}(\boldsymbol{\theta})=f(\boldsymbol{\theta}^{\star})\quad\quad\forall\,\boldsymbol{\theta},\boldsymbol{\theta}^{\star}\in\mathbb{R}^{p}\,.$ (7)
Locally identifiable parameter combinations are defined analogously [17, 18]. Subsequently, we present the standard differential algebra approach for deducing global structural identifiability.
2.2 The standard differential algebra approach
he standard differential algebra approach enables a global structural identifiability analysis for systems of ordinary differential equations whose right-hand sides $\mathbf{f}$ and output functions $\mathbf{h}$ are rational, as in Eq. (3). For such models, the approach identifies globally structurally identifiable parameter combinations through a sequence of algebraic steps [16]. First, the original state–output system in Eq. (3) is transformed into an equivalent, potentially higher-order, output-only system of differential equations, as in Eq. (4). Second, this input–output representation is rewritten as a system of monic polynomial differential equations. Third, the coefficients of these polynomials are extracted. Finally, the globally structurally identifiable parameter combinations are inferred from these coefficients.
The validity of this procedure follows from the output-equivalence of the original state–output system and the derived output system. Specifically, for any choice of parameters $\boldsymbol{\theta}$, the outputs $\mathbf{y}$ generated by a solution $\mathbf{x}$ of the original system in Eq. (3) also satisfy the corresponding output system in Eq. (4[19]. In this sense, the output system provides an exhaustive characterisation of the input–output behaviour of the original model [19]. Since solutions of ordinary differential equations are uniquely determined by their governing equations and initial conditions, the parameter combinations appearing in the output system uniquely determine the observed outputs $\mathbf{y}$, and hence correspond precisely to the globally structurally identifiable quantities.
3 Local structural identifiability and parameter symmetries
To establish the missing link between structural identifiability analyses and Lie symmetries, we first introduce the notion of parameter symmetries. Using this notion, we show that the universal parameter invariants of these parameter symmetries correspond to the locally structurally identifiable parameter quantities of a given model. Using this result, we demonstrate that the standard differential algebra approach for analysing global structural identifiability is consistent with the concept of universal parameter invariants associated with parameter symmetries.
@@ 3 Local structural identifiability and parameter symmetries
3.1 Parameter symmetries are Lie transformations acting as re-parametrisations that preserve observed outputs
Structural identifiability is defined by the requirement that changes in parameter values induce changes in the observed outputs. It is therefore natural to study transformations of the model parameters that leave the outputs invariant. Among such transformations, Lie symmetries play a distinguished role: they are infinitesimal transformations that act locally on parameters and states while preserving the model equations. Motivated by this observation, we introduce a particular class of Lie symmetries, termed parameter symmetries.
Defn. 4 (Parameter symmetries).
@@
Let $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}:\mathbb{R}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\mapsto\mathbb{R}\times\mathbb{R}^{m}\times\mathbb{R}^{p}$ be a one-parameter $\mathcal{C}^{\infty}$ diffeomorphism that is restricted to the parameters of the system of output ODEs in Eq. (4) defined by
@@
where the target functions $\hat{\boldsymbol{\theta}}$ depend solely on the parameters $\boldsymbol{\theta}$ in addition to the parameter $\varepsilon$. To simplify the notation, we simply write $\hat{\boldsymbol{\theta}}(\varepsilon)$ as opposed to $\hat{\boldsymbol{\theta}}(\boldsymbol{\theta};\varepsilon)$. The independent variable $t$ and the dependent variables $\mathbf{y}$ are invariant under the action of $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$, implying that for any solution curves the following conservation property holds:
@@
$X_{\boldsymbol{\theta}}=\sum_{\ell=1}^{p}\chi_{\ell}(\boldsymbol{\theta})\partial_{\theta_{\ell}},$ (10)
where the functions $\boldsymbol{\chi}(\boldsymbol{\theta})=\left(\chi_{1}(\boldsymbol{\theta}),\ldots,\chi_{p}(\boldsymbol{\theta})\right)$ are called the parameter infinitesimals. Then $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$ in Eq. (8) is a parameter symmetry of the system of output ODEs in Eq. (4) if its infinitesimal generator $X_{\boldsymbol{\theta}}$ solves the linearised symmetry conditions given by
@@
Essentially, a parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ is a re-parametrisation of the output system in Eq. (4) which preserves the observed outputs $\mathbf{y}$. Since these parameter symmetries are restricted to the parameters of the model, we will use the notation $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}:\boldsymbol{\theta}\mapsto\hat{\boldsymbol{\theta}}(\varepsilon)$ to describe them henceforth. Next, we define the notion of differential invariants of parameter symmetries.
Defn. 5 (Differential invariants of parameter symmetries).
Consider a parameter symmetry $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$ and its corresponding infinitesimal generator $X_{\boldsymbol{\theta}}$. A non-constant function
$I=I\left(t,\frac{\mathrm{d}^{N}\mathbf{y}}{\mathrm{d}t^{N}},\frac{\mathrm{d}^{N-1}\mathbf{y}}{\mathrm{d}t^{N-1}},\ldots,\frac{\mathrm{d}\mathbf{y}}{\mathrm{d}t},\mathbf{y},\boldsymbol{\theta}\right),$
is called a differential invariant of $\Gamma^{\boldsymbol{\theta}}_{\varepsilon}$ if it satisfies
@@
From this definition, we immediately see that the independent time variable $t$, all outputs $\mathbf{y}$, and their respective derivatives, are themselves differential invariants. In addition to the independent and dependent variables, there are other invariants solely depending on the parameters, and to distinguish between these two types of invariants we introduce the notion of parameter invariants.
Defn. 6 (Parameter invariants of parameter symmetries).
@@
Moreover, for a general output system of ODEs there are potentially many possible re-parametrisations, or, differently put, many possible parameter symmetries $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ (Defn. 11). For instance, consider the trivial parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}:\boldsymbol{\theta}\mapsto\boldsymbol{\theta}$ which is common to all possible output systems. For the trivial parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}$, every single parameter is a parameter invariant, and this is the only parameter symmetry for which this is true. Thus, for an output system with an additional parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ other than the trivial parameter symmetry, some of the parameter invariants of $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}$ are shared with $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ while other parameter invariants of the trivial parameter symmetry are unique to $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}$. In general, parameter symmetries have some parameter invariants that are unique to them, and other parameter invariants that are shared with all other parameter symmetries. We refer to the latter type of parameter invariant as universal parameter invariants.
Defn. 7 (Universal parameter invariants of a model).
@@
Importantly, the independent variable $t$, the outputs $\mathbf{y}$, and all derivatives of the outputs, are universal invariants of all parameter symmetries.
Given the system of output ODEs in Eq. (4), we know that the number of parameter invariants for any parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta}}$ is an integer in the set $\{0,\ldots,p\}$ where $p$ is the number of parameters. When we have zero parameter invariants, there are no parameters in the output ODEs, and in the case of $p$ parameter invariants, then each parameter is itself an invariant—which is only true for the trivial parameter symmetry $\Gamma_{\varepsilon}^{\boldsymbol{\theta},0}:\boldsymbol{\theta}\mapsto\boldsymbol{\theta}$. From now on, we restrict ourselves to output ODEs in the form of Eq. (4) that contain parameters, i.e. we exclude the extreme case where we have $p=0$ parameters. In particular, if the parameter $\theta_{\ell}$ for some $\ell\in\{1,\ldots,p\}$ is a parameter invariant, i.e. $I_{\boldsymbol{\theta}}(\boldsymbol{\theta})=\theta_{\ell}$, then Eq. (13) gives
@@
Consequently, a parameter $\theta_{\ell}$ is invariant if its infinitesimal is zero, i.e. $\chi_{\ell}=0$. Moreover, by generating the corresponding transformation $\hat{\theta}_{\ell}(\varepsilon)$ using the infinitesimal $\chi_{\ell}$, which entails solving the following ODE: